OpenStudy (anonymous):

A certain arithmetic sequence has the recursive formula an = an-1 + d. If the common difference between the terms of the sequence is -13, what term follows the term that has the value 13?

OpenStudy (anonymous):

\[\bf a_n=a_{n-1}+d \rightarrow a_n =13-13=?\]

OpenStudy (anonymous):

@myesha12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what's the answer? I typed everything up. @myesha12

OpenStudy (anonymous):

What goes in the place of the question mark

OpenStudy (anonymous):

i don't know that's what i was tying to figure out

OpenStudy (anonymous):

Ok basically the recursive sequence here tells us that the next term in the sequence is the previous term plus the common difference. Here we are finding the term after 13; \(\bf a_n\). So to find the term \(\bf a_n\) which comes right after 13, by the recursive sequence, we have to take the term that comes before \(\bf a_n\) and that term is 13. And we must add the common difference \(\bf d=-13\) to 13 to get \(\bf a_n\). So \(\bf 13+(-13)=13-13=0\). Get it? @myesha12

OpenStudy (anonymous):

yes!!! THANK YOU